OpenStudy (anonymous):

can someone help please How can one 1/2x-5=1/3x+6 be set up as a system of equations?

Vocaloid (vocaloid):

I'm guessing they want you to separate the two equations like so: y = 1/2x-5 y = 1/3x + 6 we know that the two functions are equal, so we can use a second variable y as an "intermediate"